OpenStudy (alison123):

how do you solve 2/3 x ^2 + x= 4 + 2/3

OpenStudy (anonymous):

i would start with \[2x^2+3x=12+2=14\]

OpenStudy (anonymous):

then \[(x-2) (2 x+7) = 0\]and you are basically done